What is the probability of randomly selecting the following sequence of boys and girls: BBGBGGB

Select one:

a. (1/2)^4 + (1/2)^3

b. (1/2)^7

c. 7*(1/2)

d. 1/2
Answer #1

Let the probablity of selecting a boy = 1/2

So, the probabilty of selecting a girl = 1/2

So for a sequence BBGBGGB

probability = (1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)

provbavility = (1/2)^7

option B is correct